Supposing a bridge is built from pieces of 4.46m long newly fabricated metal of linear expansivity 9.5×10^-6. If the gaps between the metals are 5.18mm on a day with temperature 15C, at what temperature will the pieces of metal begin to touch one another?
1
Expert's answer
2020-03-05T09:36:39-0500

Solution. By definition of the temperature coefficient of linear expansion


"\alpha = \frac {\Delta l}{l_0 \Delta t}"

where α=9.5×10^-6 0C-1 is temperature coefficient of linear expansion; ∆l=5.18 m-3; l0 =4.46 m is particle length at initial temperature t0=150C; ∆t=(t-t0) is temperature difference. Therefore, the final temperature is equal to


"t=t_0+\frac{\Delta l}{l_0 \alpha}=15+\frac{5.18\times 10^{-3}}{4.46 \times 9.5\times 10^{-6}}\approx 137^0C"

Answer.


"t= 137^0C"